Brian buys a computer for £3100.
It depreciates at a rate of 5% per year.
How much will it be worth in 5 years?
Give your answer to the nearest penny where appropriate.
please can you help me (I am stuck)

Answer by Theo(13342)

the computer costs 3100.

every year it loses 5% of its value.

how much will it be worth in 5 years.

if it loses 5% each year, then each succeeding year it is worth 95% of what it was the year before.

therefore, after 1 year is it worth .95 * what it was when you bought it.
the year after that it is worth .95 * .95.
the year after that it is worth .95 * .95 * .95.

at the end of the fifth year it is worth .95 * .95 * .95 * .95 * .95 of what it was when you bought it.

.95 ^ 5 * 3100 = 2398.720906.

the formula to use for this would be f = p * (1 + r) ^ n

f is the future value.
p is the present value.
r is the interest rate per time period (years in this case).
n is the number of time periods (years in this case).

when p = 3100 and r = -.05 and n = 5, the formula becomes:

f = 3100 * (1 - .05) ^ 5 which becomes f = 3100 * .95 ^ 5 which becomes f = 2398.720906.

to the nearest penny, the answer is 2398.72.